Perform hypothesis tests and confidence intervals for one coefficient
Answer #1

Hypothesis test for one coefficient

For the null hypothesis , assume we take \beta_{i} as the true coefficient on ith regressor which takes value \beta_{i,0}. This would form the null hypothesis . Here for generality , we are assuming a two tailed test

Null hypothesis = H0 :  \beta_{i} =  \beta_{i,0}.

Alternative hypothesis = Ha: \beta_{i}\neq \beta_{i,0}

Steps to perform hypothesis test

1) Define null and alternative hypothesiis which is done above

2) Determine the level of significance to be used

3) Calculate t statistic by first finding the standard error of the coefficient

The formula for t statistic is

z= (\bar{X}-\mu )/(\sigma /\sqrt{n})

where n is the number of obervations

\sigma /\sqrt{n} is the standard error of mean

4) Determine critical values.

5) Find whether to reject or not reject null hypothesis

6)State the conclusion

The confidence interval for one coefficient

For 95% confidence interval

95% CI= P(μ-1.96 * ơNn < X < μ 1.96 * σ/Vn) = 0.95

Rearranging it for solving for we get

P(X-1.96 *o/Vn < μ < X + 1.96 *o/v/r) = 0.95

So the confidence interval would be

(X-1.96 * σ/v/r. X + 1.96 * σ/vn)